Circular No. 6963

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1E 1740.7–2942 D. M. Smith, Space Sciences Laboratory, University of California at

Berkeley (UCB/SSL); S.-N. Zhang, Universities Space Research Association and Marshall Space Flight Center (MSFC), NASA; D. Main, UCB/SSL; and B. A. Harmon, MSFC, report: “This black-hole candidate and microquasar is entering a hard-x-ray low-flux state for the first time in three years. The decline was measured by CGRO/BATSE beginning between Mar. 18 and Apr. 7, from between 100 and 150 mCrab (20–100 keV) to ∼ 30–50 percent of that level at present. In the RXTE PCA (2.5–25 keV), a spectral softening began around Mar. 21 from power-law index 1.5 to 1.7, but the count rate only declined after Apr. 20 and is now at 60 percent of maximum. Comparison with similar declines in 1990–1991 and 1994 (BATSE and SIGMA data) indicates a possible period of 1300–1400 days. If the pattern repeats, the low state should last several months. Observations are strongly encouraged at all wavelengths not excessively absorbed near the Galactic center.”

NOVA SAGITTARII 1998 Y. Lipkin, A. Retter, and E. M. Leibowitz, Wise Observatory, Tel Aviv

University, report: “On June 7, 8, 9, 11, 12, and 13, we performed CCD photometry of N Sgr 98 with the 1-m Wise telescope through an I filter for ∼ 5 hr each night. The light curve is modulated coherently for at least 40 cycles, with either one of the two periodicities (0.17330 or 0.14765 ± 0.00011 day), which are 1-day aliases of each other. The folded data resemble an eclipse light curve, with an average dip of 0.014 mag. It is likely that similar periodicity modulates the light curve of the star as measured by us already on May 11 and 12.”

Visual magnitude estimates by K. Hornoch, Lelekovice, Czech Republic: June 20.96 UT, 11.3; 21.96, 11.3; 26.00, 11.4; 28.96, 11.3.

COMET C/1998 J1 (SOHO) Visual m_{1} estimates: June 8.90 UT, 6.1 (H. C. Vital, Rio de Janeiro,

Brazil, 12×50 binoculars); 12.94, 6.8 (W. C. de Souza, São Paulo, Brazil, 11×80 binoculars); 18.32, 7.0 (S. T. Rae, Palmerston North, N.Z., 10×50 binoculars); 23.91, 7.6 (de Souza); 28.28, 7.6 (Rae, Te Awamutu, N.Z.).

1998 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green